# An algorithm for estimating the crossing number of dense graphs, and continuous analogs of the crossing and rectilinear crossing numbers

### v2

@@ 3 Estimating the crossing number
Figure 2: A section of $C_{i}$ (red) is contained in the boundary of $R_{i}$ and the two adjacent arcs (also in red) lie inside $R_{i}$. For every vertex in this section, consider the edges of $G$ that are incident to it, lie in $R_{i}$, and are contained in a curve that represents an edge of $G$ in $\mathcal{D}$. Then, modify $C_{i}$ as shown above. Since no vertex of $G$ belongs to $C_{i}$, this increases the number of nodes in the curve by at most three times the number of nodes that conform the original section of $C_{i}$.
@@
It is not hard to see that the number of fundamental regions is $O(1/\varepsilon^{2})$, but we will also require an upper bound on the total number of arcs that conform their boundaries. This quantity is at most the total number of nodes of the $C_{i}^{\prime}$’s. The argument that follows is similar to the one used in the proof of Corollary 5 from [34]. The level of a region $R_{i}$, which we denote by $l(R_{i})$, is defined as follows: If $R_{i}$ is fundamental, then its level is 00. Otherwise, its level is obtained by adding $1$ to the largest of the levels of the two regions that were created when splitting $R_{i}$ using $C_{i}^{\prime}$. The key observation here is that any two $R_{i}$’s with the same level are interior disjoint. This can be seen to imply that the orders of the $T_{i}$’s such that $l(R_{i})=m$ for a fixed $m$ add up to
$O\left(n^{4}+\sum_{t,\ l(R_{t})<m}|C^{\prime}_{t}|\right).$
@@
But it is not hard to check that the total complexity of the $C_{t}^{\prime}$’s is $o(n^{4})$, so the above expression is $O(n^{4})$. On the other hand, every $R_{i}$ with level $m\geq 1$ contains at least $\lceil\varepsilon^{2}n\rceil(3/2)^{m-1}$ vertices of $G$, so there are at most $\varepsilon^{-2}(2/3)^{m-1}$ such regions. Now, the Cauchy-Schwartz inequality yields that the total complexity of the $C_{i}^{\prime}$’s in level $m$ is at most
@@
$\sum_{i,\ l(R_{i})=m}4(L\sqrt{m_{i}}+3n_{i})\leq L^{\prime}\frac{n^{2}}{\varepsilon}\left(\frac{2}{3}\right)^{\frac{m-1}{2}}+12n^{2},$
@@
$12n^{2}\cdot l(R_{0})+\sum_{m=1}^{l(R_{0})}L^{\prime}\frac{n^{2}}{\varepsilon}\sqrt{\frac{2}{3}}^{(m-1)}=O(n^{2}/\varepsilon),$
since $l(R_{0})=O(\log\varepsilon^{-1})$. We remark that the bound on the number of regions with level $m$ also implies that there are $O(1/\varepsilon^{2})$ regions in total.
Next, we show how the fundamental regions can be further subdivided into smaller parts that satisfy the requirements in the statement of the theorem. Classify the edges of $G$ as light or heavy depending on whether or not they are involved in less than $\varepsilon n^{2}/16$ crossings, respectively. By applying the crossing lemma to the subgraph that contains only the light edges, we get that the number of such edges is $O(\varepsilon^{1/2}n^{2})$. For any two points $p$ and $q$ on the plane, define $d_{\mathcal{D}}(p,q)$ as follows: If $p$ and $q$ belong to the same fundamental region and at least one of them does not lie on its boundary, then let $d_{\mathcal{D}}$ be the least positive integer such that there exists a clean curve which has endpoints $p$ and $q$, is contained in that region, and whose interior has at most that many intersection points with the edges of the drawing. Otherwise, set $d_{\mathcal{D}}(p,q)=\infty$. For any point $p$ on the plane and any positive integer $k$, let $B_{\mathcal{D}}(p,k)$ be the set that consists of those points $q$ such that $d_{\mathcal{D}}(p,q)\leq k$, and note that it is connected. We say that an edge is encapsulated by $B_{\mathcal{D}}(p,k)$ if both of its endpoints and all the points where it crosses another edge are contained in $B_{\mathcal{D}}(p,k)$, and at least one of these points belongs to the interior of this region.
@@
For every point $p$ that does not belong to the boundary of a fundamental region, $B_{\mathcal{D}}(p,\varepsilon n^{2}/4)$ (which we assume is not the whole plane) satisfies at least one of the properties below.
@@
    It contains at least $\varepsilon n^{2}/16$ arcs of the boundary of the fundamental region that encloses it.
@@
    It contains one of the $C_{i}^{\prime}$’s in its entirety.
@@
First, we deal with the case where $B_{\mathcal{D}}(p,3\varepsilon n^{2}/16)$ contains a point from the boundary of the corresponding fundamental region. By starting from this point and walking along the said boundary in both directions, we will either encounter at least $\varepsilon n^{2}/16$ arcs that are contained in $B_{\mathcal{D}}(p,\varepsilon n^{2}/4)$ (the fact that no vertex of $G$ is contained in the boundary of a fundamental region is important here), or go around the entirety of one of the $C_{i}^{\prime}$’s. Hence at least one of the first two properties holds.
@@
Suppose now that $B_{\mathcal{D}}(p,\varepsilon n^{2}/16)$ contains a heavy edge $e$. Then, we can find at least $\varepsilon n^{2}/16$ crossing points of $\mathcal{D}$ that involve $e$ and are contained in $B_{\mathcal{D}}(p,\varepsilon n^{2}/8)$; apart from $e$, the edges involved in these crossings are all distinct due to the fact that $\mathcal{D}$ is simple. If any of these edges is light, then it must be encapsulated by $B_{\mathcal{D}}(p,3\varepsilon n^{2}/16)$, so we can assume that at least $\varepsilon n^{2}/32$ of them are heavy. Each of these heavy edges takes part in at least $\varepsilon n^{2}/16$ crossing that lie in $B_{\mathcal{D}}(p,3\varepsilon n^{2}/16)$. Summing over all of these edges and dividing by two to account for the fact that each crossing might be counted up to two times, we get that at least $(\varepsilon n^{2}/32)(\varepsilon n^{2}/16)/2$ crossing points of $\mathcal{D}$ belong to the interior of $B_{\mathcal{D}}(p,\varepsilon n^{2}/4)$.
@@
It remains to tackle the case where the interior of $B_{\mathcal{D}}(p,\varepsilon n^{2}/16)$ is disjoint from all the heavy edges. Given a clean curve $C$, let $c_{1},c_{2},\dots c_{t}$ denote the intersection points of its interior with the edges of $G$, and let $e_{i}$ be the edge that contains $c_{i}$; we define $w_{G}(C)$ as $\sum_{i=1}^{t}w_{G}(e_{i})$. Consider a clean curve $C_{p}$ that connects $p$ to a point on the boundary of $B_{\mathcal{D}}(p,\varepsilon n^{2}/16)$ which is neither a crossing nor a vertex of $G$, and such that $w(C_{p})$ is minimal amongst all curves of this kind. Furthermore, suppose that there is no curve with the aforementioned properties which has less intersection points with the edges than $C_{p}$ does. Each of the at least $\varepsilon n^{2}/16$ points where the interior of $C_{p}$ intersects an edge must belong to a light edge. Since $\mathcal{D}$ is locally optimal, no two of these points belong to the same edge, as otherwise we would be able to construct a curve that contradicts the choice of $C_{p}$. Hence there are at least $\varepsilon n^{2}/16$ light edges that intersect the interior of $B_{\mathcal{D}}(p,\varepsilon n^{2}/16)$ and are thus encapsulated by $B_{\mathcal{D}}(p,\varepsilon n^{2}/4)$. ∎
@@
The desired subdivision can now be obtained by means of a standard covering argument. Indeed, let $P=\{p_{1},p_{2},\dots,p_{k}\}$ be a set of points on the plane such that none of its elements represents a vertex or lies on an edge or the boundary of a fundamental region, and which is maximal with the property that $B_{\mathcal{D}}(p_{i},\varepsilon n^{2}/4)$ and $B_{\mathcal{D}}(p_{j},\varepsilon n^{2}/4)$ are interior disjoint whenever $i\neq j$. Given that the total complexity of the boundaries of the fundamental regions is $O(n^{2}/\varepsilon)$, the number of $B_{\mathcal{D}}(p_{i},\varepsilon n^{2}/4)$’s that satisfy property $1$ is $O(1/\varepsilon^{2})$. Since the total number $C_{i}^{\prime}$’s is $O(1/\varepsilon^{2})$, the same is true for those that satisfy property $2$. Because the $B_{\mathcal{D}}(p_{i},\varepsilon n^{2}/4)$’s are interior disjoint, every crossing point of $\mathcal{D}$ belongs to the interior of at most one of them, and no edge can be encapsulated by more than one of them. Since $\mathcal{D}$ is simple, the number of crossing points is $O(n^{4})$, so at most $O(1/\varepsilon^{2})$ of the $B_{\mathcal{D}}(p_{i},\varepsilon n^{2}/4)$’s contain $\varepsilon^{2}n^{4}/32^{2}$ such points or more. Lastly, as there are no more than $O(\varepsilon^{1/2}n^{2})$ light edges, the amount of $B_{\mathcal{D}}(p_{i},\varepsilon n^{2}/4)$’s that encapsulate at least at $\varepsilon n^{2}/32$ light edges is upper bounded by $O(1/\varepsilon^{2})$. This shows that $k=O(1/\varepsilon^{2})$. Now, notice that for every point $q$ we can find a $p_{i}$ such that such that $d_{\mathcal{D}}(p_{i},q)\leq\varepsilon n^{2}/2$, or else $q$ could be added to $P$, contradicting its maximality. For each $q$, let $P(q)\subseteq P$ denote the set of $p_{i}$’s that minimize $d_{\mathcal{D}}(p_{i},q)$, and let $p(q)$ be the element of $p(q)$ with the smallest subscript. Define $r_{i}$ as the closure of $\{q\ |\ p(q)=p_{i}\}$ for every $i\in[k]$ [^7] If $m$ is a positive integer, then $[m]$ denotes the set $\{1,2,\dots,m\}$.. The $r_{i}$’s form a subdivision of the plane, and they are closed, connected and interior disjoint. Moreover, if $q\in r_{i}$ and $C_{q}$ is a clean curve that connects $q$ to $r_{i}$ such that its interior has precisely $d_{\mathcal{D}}(p_{i},p)$ intersection points with the edges of $G$, then it is not hard to see that this curve lies completely within $r_{i}$. It follows that any two points in $r_{i}$ can be connected by a clean curve that goes through $p_{i}$ and whose interior crosses at most $\varepsilon n^{2}$ edges. Lastly, the $r_{i}$’s can be modified using the process depicted in Figure 1 so as to ensure that no vertex of $G$ lies on the boundary of any of them (this final step is where specific wording of property III comes into play). The resulting regions satisfy properties I, II and III.
As promised, we now show that there is a strong connection between $d_{\square}$ and the crossing number.
@@
Let $G_{1}$ and $G_{2}$ be edge weighted graphs on the same vertex set $V$ and let $d=d_{\square}(G_{1},G_{2})$. If $d\geq n^{-4}$, then
@@
where $M$ is an absolute constant. Meanwhile, $|\mathop{\mathrm{cr}}(G_{1})-\mathop{\mathrm{cr}}(G_{2})|\leq d^{2}n^{8}$ holds independently of the value of $d$.
Remark: The interesting part of this result is the first inequality.
@@
The proof of Theorem 1.1 by Fox, Pach and Suk relies on a regularity lemma for semi-algebraic graphs from*. In a sense, Theorem 3.1 will act as a purely combinatorial substitute of the said lemma, although this analogy should not be pushed too far.
@@
Suppose that $d\geq n^{-4}$. By adding edges of weight 00, we can assume that every two elements of $V$ are adjacent in $G_{1}$. Consider a simple and locally optimal drawing $\mathcal{D}$ of $G_{1}$ and apply Theorem 3.1 with $\varepsilon=d^{1/8}$ to obtain a subdivision of the plane. Let $r_{1},r_{2},\dots,r_{k}$ be the regions that contain at least one vertex and, for each $i$, let $V_{i}\subset V$ denote the set of vertices that are contained in $r_{i}$. We can assume that $k\leq C/\varepsilon^{2}=C/d^{1/4}$ for some absolute constant $C$. Since $\varepsilon\geq n^{-1/2}$, each of the $V_{i}$’ has at most $\lceil\varepsilon^{2}n\rceil<2\varepsilon^{2}n$ elements. We say that an edge is long if its endpoint belong to different $V_{i}$’s, and that it is short otherwise.
@@
We shall use $\mathcal{D}$ as a basis for constructing a drawing $\mathcal{D}^{\prime}$ of $G_{2}$. The elements of $V$ will be represented by the same points as in $\mathcal{D}$. Next, for every long edge $(u,v)\in E_{G_{2}}(V_{i},V_{j})$, choose a random edge $(u^{\prime},v^{\prime})\in E_{G_{1}}(V_{i},V_{j})$, where $(x,y)$ is chosen with probability $w_{G_{1}}(x,y)/e_{G_{1}}(V_{i},V_{j})$ (these selections are carried out independently of each other for every long edge of $G_{2}$). If $e_{G_{1}}(V_{i},V_{j})=0$, $(u^{\prime},v^{\prime})$ is chosen uniformly at random from $E_{G_{1}}(V_{i},V_{j})$. Consider the points where the curve that represents $(u^{\prime},v^{\prime})$ in $\mathcal{D}$ crosses the union of the boundaries of $r_{i}$ and $r_{j}$ (if a section of the curve is contained in the union of the boundaries, we take only its endpoints) and number them by $p_{1},p_{2},\dots,p_{t}$ in the order that they appear as this curve is traversed from $u$ to $v$. If some $p_{t}$ lies on the boundaries of both $r_{i}$ and $r_{j}$, then write $a(u,v)=b(u,v)=p_{t}$. Otherwise, choose two contiguous $p_{i}$’s $a(u,v)=p_{t}$ and $b(u,v)=p_{t\pm 1}$ such that $a(u,v)$ lies on the boundary of $r_{i}$ and $b(u,v)$ lies on the boundary of $r_{j}$. The curve that represents $(u,v)$ will be composed of three sections: one which goes from $u$ to $a(u,v)$, one that goes from $a(u,v)$ to $b(u,v)$ by following along the curve that represents $(u^{\prime},v^{\prime})$ in $\mathcal{D}$ (which is a single point in the case that $a(u,v)=b(u,v)$), and a last one between $b(u,v)$ and $v$. The first section will be constructed so that it is fully contained in $r_{i}$ and it has as few crossings as possible with the edges of $G_{1}$ that have no endpoint in $V_{i}$. Analogously, we draw the last section so that it lies in $r_{j}$ and has the least possible number of crossings with the edges of $G_{1}$ that have no endpoint in $V_{j}$. The first and last sections of every curve that represents a long edge of $G_{2}$ will be called extremal. Each short edge with both endpoints in some $V_{i}$ will be represented by a curve that is contained in $r_{i}$ and has the least possible number of crossings with the the edges of $G_{1}$ that have no endpoint in $V_{i}$. See Figure 3. Let $E_{i}$ denote the set of curves formed by the short edges of $G_{2}$ with endpoints in $V_{i}$, and the extremal sections with an endpoint in $V_{i}$. The above construction can be carried out so that each element of $E_{i}$ has at most $\varepsilon n^{2}$ crossings with the edges of $G_{1}$ that have no endpoint in $V_{i}$, and such that any two elements of $E_{i}$ have a finite number of points in common. There might still be some pairs of elements of $E_{i}$ which cross each other more than once, but these multiple crossings can be eliminated without increasing the total number of crossing between the elements of $E_{i}$ and the edges of $G_{1}$ with no endpoint in $V_{i}$. Finally, observe that some pairs of long edges of $E_{G_{2}}(V_{i},V_{j})$ may have infinitely many points in common if their non extremal sections coincide. Thus, in order to obtain an actual graph drawing we might have to perturb these edges slightly as in Figure 4. We point out that the resulting drawing is not necessarily simple.
@@
Figure 3: On the right, we have the curve that represents $(u^{\prime},v^{\prime})$ in $\mathcal{D}$. Setting $a(u,v)=p_{4}$ and $b(u,v)=p_{5}$ and proceeding as described in the proof, we arrive at a curve that connects $u$ and $v$.

### v4

@@ 3.1 Subdividing drawings
Figure 2: The region $R_{i}$ is shaded. A section of $C_{i}^{\prime}$ (red) is contained in the boundary of $R_{i}$, and the two adjacent arcs of $C_{i}^{\prime}$ (also in red) lie inside $R_{i}$. For every vertex of $T_{i}^{\prime}$ in this section, consider the edges of $T_{i}^{\prime}$ that are incident to it, lie in $R_{i}$, and are contained in a curve that represents an edge of $G$ in $\mathcal{D}$. Note that, since no vertex of $G$ belongs to $C_{i}$, there are at most two edges of $G$ going through any of these vertices. Now, modify $C_{i}^{\prime}$ as shown above. This increases the number of nodes in the curve by at most three times the number of nodes that conform the original section of $C_{i}$.
After enough steps, we reach a subdivision of the plane into closed, connected and interior disjoint regions, each of which contains at most $\lceil\varepsilon^{2}n\rceil$ vertices of $G$. These regions will be referred to as fundamental regions from now on. An important feature of above procedure that will prove useful later on is that each connected component of the boundary of a region $R_{j}$ contains one of the $C_{i}^{\prime\prime}$’s in its entirety (possibly, but not necessarily, $C_{j}^{\prime\prime}$ itself). The number of fundamental regions is clearly $O(1/\varepsilon^{2})$, but we also require an upper bound on the total number of intersections between their boundaries and the edges of $G$ in the drawing. A quick inspection of the subdivision process reveals that this quantity is at most twice the sum of the numbers of nodes in $C_{0}^{\prime}$ and in each of the $C_{i}^{\prime\prime}$’s. The counting argument that is used below to bound this quantity is similar to the one from the proof of Corollary 5 in [40].
@@
The level of a region $R_{i}$, which we denote by $\ell(R_{i})$, is defined as follows: If $R_{i}$ is fundamental, then its level is 00. Otherwise, its level is obtained by adding $1$ to the largest of the levels of the two regions that were created when splitting $R_{i}$ using $C_{i}^{\prime\prime}$. The key observation here is that any two $R_{i}$’s with the same level are interior disjoint. This implies that the orders $m_{i}$ of the $T_{i}$’s such that $\ell(R_{i})=l$ for a fixed $l$ add up to no more than
$O\left(n^{4}+|C^{\prime}_{0}|+\sum_{i>0\text{ with }\ell(R_{i})>l}|C^{\prime\prime}_{i}|\right).$
@@
The largest level (which is $\ell(R_{0})$) is clearly no more than $O(\log n)$, so one can check inductively that the total complexity of the $C_{i}^{\prime\prime}$’s with level $l$ is by $o(n^{3})$, so the above expression is $O(n^{4})$ Also, every $R_{i}$ with level $l\geq 1$ contains at least $\lceil\varepsilon^{2}n\rceil(3/2)^{l-1}$ vertices of $G$, so there are at most $\varepsilon^{-2}(2/3)^{l-1}$ such regions. Now, the Cauchy-Schwarz inequality yields that the total complexity of the $C_{i}^{\prime\prime}$’s corresponding to regions in level $l$ is bounded from above by
@@
$\sum_{i\text{ with }\ell(R_{i})=l}4(L\sqrt{19m_{i}}+3n_{i})\leq L^{\prime}\frac{n^{2}}{\varepsilon}\left(\frac{2}{3}\right)^{\frac{l-1}{2}}+12n,$
@@
$12n\cdot\ell(R_{0})+L^{\prime}\frac{n^{2}}{\varepsilon}\sum_{l=1}^{\ell(R_{0})}\sqrt{\frac{2}{3}}^{(l-1)}=O(n^{2}/\varepsilon),$
where we have used the fact that $\ell(R_{0})=O(\log\varepsilon^{-1})=o(n)$. Note that the bound on the number of regions with level $l$ also yields that there are $O(1/\varepsilon^{2})$ regions in total, not just at level 00. Indeed, there are at most
$\sum_{l=1}^{\ell(R_{0})}\varepsilon^{-2}(2/3)^{l-1}=O(1/\epsilon^{2})$
regions with level $l\geq 1$. This observation will be used later.
Next, we show how the fundamental regions can be further subdivided into smaller parts that satisfy the requirements in the statement of the theorem. Let us start by classifying the edges of $G$ as light or heavy depending on whether or not they are involved in less than $\varepsilon n^{2}/16$ crossings, respectively. By applying the crossing lemma to the subgraph that contains only the light edges, we get that the number of such edges is $O(\varepsilon^{1/2}n^{2})$. A simple curve whose relative interior does not pass through any vertex or crossing point of $\mathcal{D}$ will be called clean. Now, for any two points $p$ and $q$ on the plane, we define the distance $d_{\mathcal{D}}(p,q)$ as follows:
If $p$ and $q$ belong to the same fundamental region and at least one of them does not lie on its boundary, then $d_{\mathcal{D}}$ is the least positive integer $k$ such that there exists a clean curve which has endpoints $p$ and $q$, is contained in that region, and whose interior has at most $k$ intersection points with the edges of the drawing. Otherwise, set $d_{\mathcal{D}}(p,q)=\infty$.
For any point $p$ on the plane and any positive number $k$, let $B_{\mathcal{D}}(p,k)$ be the set that consists of those points $q$ such that $d_{\mathcal{D}}(p,q)\leq k$, and note that $B_{\mathcal{D}}(p,k)$ is closed and connected. We say that an edge of $G$ is encapsulated by $B_{\mathcal{D}}(p,k)$ if both of its endpoints and all the points where it crosses another edge are contained in $B_{\mathcal{D}}(p,k)$, and at least one of these points belongs to the interior of this region.
@@
For every point $p$ that does not belong to the boundary of a fundamental region, if $B_{\mathcal{D}}(p,\varepsilon n^{2}/4)$ is not the whole plane then it satisfies at least one of the properties below.
@@
    It contains at least $\varepsilon n^{2}/16$ intersections between the boundary of the fundamental region that contains $p$ and the edges of the drawing.
@@
    It contains one of the $C_{i}^{\prime\prime}$’s in its entirety.
@@
First, we deal with the case where $B_{\mathcal{D}}(p,3\varepsilon n^{2}/16)$ contains a point from the boundary of the fundamental region $R_{t}$ that contains $p$. By starting from this point and walking along the corresponding connected component of the boundary of $R_{t}$ in both directions, we will either encounter at least $\varepsilon n^{2}/16$ intersections between the edges and the boundary of the region, all of which lie in $B_{\mathcal{D}}(p,\varepsilon n^{2}/4)$ (the fact that no vertex of $G$ is contained in the boundary of a fundamental region is important here), or will go around the entirety of one of the $C_{i}^{\prime\prime}$’s. Hence, at least one of the first two properties holds. From now on, we assume that $B_{\mathcal{D}}(p,3\varepsilon n^{2}/16)$ does not contain a point from the boundary of the corresponding fundamental region.
@@
Suppose now that $B_{\mathcal{D}}(p,\varepsilon n^{2}/16)$ contains a point from a heavy edge $e$. Then, we can find at least $\varepsilon n^{2}/16$ crossing points of $\mathcal{D}$ that involve $e$ and are contained in $B_{\mathcal{D}}(p,\varepsilon n^{2}/8)$. Apart from $e$, the edges involved in these crossings are all distinct due to the fact that $\mathcal{D}$ is simple. If any of these edges is light, then it must be encapsulated by $B_{\mathcal{D}}(p,3\varepsilon n^{2}/16)$, so we can assume that at least $\varepsilon n^{2}/32$ of them are heavy. Each of these heavy edges takes part in at least $\varepsilon n^{2}/16$ crossing that lie in $B_{\mathcal{D}}(p,3\varepsilon n^{2}/16)$. Summing over all of these edges and dividing by two to account for the fact that each crossing might be counted up to two times, we get that at least $(\varepsilon n^{2}/32)(\varepsilon n^{2}/16)/2$ crossing points of $\mathcal{D}$ belong to the interior of $B_{\mathcal{D}}(p,\varepsilon n^{2}/4)$.
@@
It remains to tackle the case where the interior of $B_{\mathcal{D}}(p,\varepsilon n^{2}/16)$ is disjoint from all the heavy edges. Given a clean curve $C$, let $c_{1},c_{2},\dots c_{t}$ denote the intersection points of its interior with the edges of $G$, and let $e_{i}$ be the edge that contains $c_{i}$; we define $w_{G}(C)$ as $\sum_{i=1}^{t}w_{G}(e_{i})$. Consider a clean curve $C_{p}$ that connects $p$ to a point on the boundary of $B_{\mathcal{D}}(p,\varepsilon n^{2}/16)$ which is neither a crossing nor a vertex of $G$, and such that $w(C_{p})$ is as small as possible amongst all curves of this kind. Furthermore, suppose that there is no curve with the aforementioned properties which has less intersection points with the edges of $G$ than $C_{p}$ does. Each of the at least $\varepsilon n^{2}/16$ points where the interior of $C_{p}$ intersects an edge must belong to a light edge. Since $\mathcal{D}$ is locally optimal, no two of these points belong to the same edge, as otherwise we would be able to construct a curve that contradicts the choice of $C_{p}$ by rerouting a portion of $C_{p}$ along such an edge. Hence, there are at least $\varepsilon n^{2}/16$ light edges that intersect the interior of $B_{\mathcal{D}}(p,\varepsilon n^{2}/16)$ and are thus encapsulated by $B_{\mathcal{D}}(p,\varepsilon n^{2}/4)$. ∎
@@
We go back to the proof of Theorem 3.1. The desired subdivision can now be obtained by means of a standard covering argument. Indeed, let $P=\{p_{1},p_{2},\dots,p_{k}\}$ be a set of points on the plane such that none of its elements represents a vertex or lies on an edge or on the boundary of a fundamental region, and which is maximal with the property that $B_{\mathcal{D}}(p_{i},\varepsilon n^{2}/4)$ and $B_{\mathcal{D}}(p_{j},\varepsilon n^{2}/4)$ are interior disjoint whenever $i\neq j$. Given that the total complexity of the boundaries of the fundamental regions is $O(n^{2}/\varepsilon)$, the number of $B_{\mathcal{D}}(p_{i},\varepsilon n^{2}/4)$’s that satisfy property $1$ is $O(1/\varepsilon^{2})$. Since the total number of $C_{i}^{\prime\prime}$’s is $O(1/\varepsilon^{2})$ and there can be at most two regions which contain a specific $C_{i}^{\prime\prime}$ in its entirety, the same is true for those that satisfy property $2$. Because the $B_{\mathcal{D}}(p_{i},\varepsilon n^{2}/4)$’s are interior disjoint, every crossing point of $\mathcal{D}$ belongs to the interior of at most one of them, and no edge can be encapsulated by more than one of them. Since $\mathcal{D}$ is simple, the number of crossing points is $O(n^{4})$, so at most $O(1/\varepsilon^{2})$ of the $B_{\mathcal{D}}(p_{i},\varepsilon n^{2}/4)$’s satisfy property 3. Lastly, as there are no more than $O(\varepsilon^{1/2}n^{2})$ light edges, the amount of $B_{\mathcal{D}}(p_{i},\varepsilon n^{2}/4)$’s that encapsulate at least at $\varepsilon n^{2}/32$ light edges is upper bounded by $O(1/\varepsilon^{2})$. This shows that $k=O(1/\varepsilon^{2})$.
Next, notice that for every point $q$ we can find a $p_{i}$ such that such that $d_{\mathcal{D}}(p_{i},q)\leq\varepsilon n^{2}/2$, or else $q$ could be added to $P$, contradicting its maximality. For each $q$, let $P(q)\subseteq P$ denote the set of $p_{i}$’s that minimize $d_{\mathcal{D}}(p_{i},q)$, and let $p(q)$ be the element of $P(q)$ with the smallest subscript. Define $r_{i}$ as the closure of $\{q\ |\ p(q)=p_{i}\}$ for every $i\in\{1,\dots,k\}$. The $r_{i}$’s form a subdivision of the plane, and they are closed, connected and interior disjoint. Moreover, if $q\in r_{i}$ and $C_{q}$ is a clean curve that connects $q$ to $p_{i}$ such that its interior has precisely $d_{\mathcal{D}}(p_{i},q)$ intersection points with the edges of $G$, then this curve lies completely within $r_{i}$. It follows that any two points in $r_{i}$ can be connected by a clean curve that goes through $p_{i}$ and whose interior crosses at most $\varepsilon n^{2}$ edges. Lastly, the $r_{i}$’s can once again be modified using a process similar to the one depicted in Figure 1 so as to ensure that no vertex of $G$ lies on the boundary of any of them (this final step is where specific wording of property (III) comes into play). The resulting regions satisfy properties (I), (II) and (III).
3.2 Cut distance and crossing numbers
The proof of Theorem 1.1 by Fox, Pach and Suk relies on a regularity lemma for semi-algebraic graphs, also by the same authors [20]. In some sense, Theorem 3.1 will act as a purely combinatorial substitute of the said lemma (although this analogy should not be pushed too far). We will now use this theorem to show that any two graphs which are close with respect to $d_{\square}$ have similar crossing numbers.
@@ 3.2 Cut distance and crossing numbers
Let $G_{1}$ and $G_{2}$ be edge-weighted $n$-vertex graphs on the same vertex set $V$ and write $d=d_{\square}(G_{1},G_{2})$. If $d\geq n^{-4}$, then
@@
where $M$ is an absolute constant. Moreover, $|\mathop{\mathrm{cr}}(G_{1})-\mathop{\mathrm{cr}}(G_{2})|\leq d^{2}n^{8}$ holds unconditionally.
We should that the content of this theorem lies in the first inequality.
@@
Suppose that $d\geq n^{-4}$. By adding edges of weight 00, we can assume that every two elements of $V$ are adjacent in $G_{1}$. Consider a simple and locally optimal drawing $\mathcal{D}$ of $G_{1}$ and apply Theorem 3.1 with $\varepsilon=d^{1/8}$ to obtain a subdivision of the plane. Let $r_{1},r_{2},\dots,r_{k}$ be the regions that contain at least one vertex and, for each $i$, let $V_{i}\subset V$ denote the set of vertices that are contained in $r_{i}$. We have that $k\leq C/\varepsilon^{2}=C/d^{1/4}$ for some absolute constant $C$. Since $\varepsilon\geq n^{-1/2}$, each of the $V_{i}$’s has at most $\lceil\varepsilon^{2}n\rceil<2\varepsilon^{2}n$ elements. We say that an edge is long if its endpoints belong to different $V_{i}$’s, and that it is short otherwise.
@@
We shall use $\mathcal{D}$ as a blueprint for constructing a drawing $\mathcal{D}^{\prime}$ of $G_{2}$. The elements of $V$ will be represented by the same points as in $\mathcal{D}$. Next, for every long edge $(u,v)\in E_{G_{2}}(V_{i},V_{j})$, choose a random edge $(u^{\prime},v^{\prime})\in E_{G_{1}}(V_{i},V_{j})$, where $(x,y)$ is chosen with probability $w_{G_{1}}(x,y)/e_{G_{1}}(V_{i},V_{j})$ (if $e_{G_{1}}(V_{i},V_{j})=0$, then $(u^{\prime},v^{\prime})$ is chosen uniformly at random from $E_{G_{1}}(V_{i},V_{j})$). These selections are carried out independently from each other for every long edge of $G_{2}$. Consider the points where the curve that represents $(u^{\prime},v^{\prime})$ in $\mathcal{D}$ crosses the union of the boundaries of $r_{i}$ and $r_{j}$ (if a section of the curve is contained in the union of the boundaries, we take only its endpoints) and label them as $p_{1},p_{2},\dots,p_{t}$ in the order that they appear as this curve is traversed from $u$ to $v$. If some $p_{i}$ lies on the intersection of the boundaries of $r_{i}$ and $r_{j}$, then we write $a(u,v)=b(u,v)=p_{i}$. Otherwise, we can choose two points $a(u,v)=p_{i}$ and $b(u,v)=p_{i\pm 1}$ such that $a(u,v)$ lies on the boundary of $r_{i}$ and $b(u,v)$ lies on the boundary of $r_{j}$. The curve that represents $(u,v)$ will be composed of three sections: one which goes from $u$ to $a(u,v)$, one that goes from $a(u,v)$ to $b(u,v)$ by following along the curve that represents $(u^{\prime},v^{\prime})$ in $\mathcal{D}$ (which is a single point in the case that $a(u,v)=b(u,v)$), and one last section from $b(u,v)$ to $v$. The first section will be constructed so that it is fully contained in $r_{i}$ and it has as few crossings as possible with the edges of $G_{1}$ that have no endpoint in $V_{i}$. Analogously, we will draw the last section so that it lies within $r_{j}$ and has the least possible number of crossings with the edges of $G_{1}$ that have no endpoint in $V_{j}$. The first and last sections of every curve that represents a long edge of $G_{2}$ will be called extremal. See Figure 3. Each short edge with both endpoints in some $V_{i}$ will be represented by a curve that is contained in $r_{i}$ and has the least possible number of crossings with the edges of $G_{1}$ that have no endpoint in $V_{i}$.
@@
Figure 3: On the right, we have the curve that represents $(u^{\prime},v^{\prime})$ in $\mathcal{D}$. Setting $a(u,v)=p_{4}$ and $b(u,v)=p_{3}$ and proceeding as described in the proof, we arrive at a curve that connects $u$ and $v$. Note that we could have also chosen $a(u,v)=p_{1}$ and $b(u,v)=p_{2}$, or $a(u,v)=p_{5}$ and $b(u,v)=p_{6}$.
Let $E_{i}$ denote the set of curves which consists of the short edges of $G_{2}$ with endpoints in $V_{i}$, and the extremal sections with an endpoint in $V_{i}$. By property (III) in the statement of Theorem 3.1, the above construction can be carried out so that each element of $E_{i}$ has at most $\varepsilon n^{2}$ crossings with the edges of $G_{1}$ that have no endpoint in $V_{i}$, and such that any two elements of $E_{i}$ have a finite number of points in common. There might still be some pairs of elements of $E_{i}$ which cross more than once, but these multiple crossings can be eliminated by rerouting some of the curves without increasing the total number of crossing between the elements of $E_{i}$ and the edges of $G_{1}$ with no endpoint in $V_{i}$. Finally, observe that some pairs of long edges of $E_{G_{2}}(V_{i},V_{j})$ may have infinitely many points in common if their non-extremal sections coincide. Thus, in order to obtain an actual graph drawing, we might have to perturb these edges slightly, as shown in Figure 4. We point out that the resulting drawing $\mathcal{D}^{\prime}$ is not necessarily simple.